How many liters of a 2% salt solution must be added to 30 liters of a 10% salt solution to dilute it to an 8% solution?
Answer by mananth(16946)

2% salt solution x liters
30 liters of a 10%
8% solution (30+x) liters
..
0.02x+30*0.1=0.08*(30+x)
0.02x +3=2.4+0.08x
Add -0.02x
0.02x-0.02x+3 = 2.4 +0.08x-0.02x
3=2.4 +0.06x
add -2.4
3-2.4= 2.4-2.4+0.06x
0.6 = 0.06x
divide by 0.06
0.6/0.06= 0.06x/0.06
x= 10 liters